Hello,
I have a word problem question I need help with.
A bank teller has some five dollar bills and some twenty dollar bills. The teller has five more twenties than fives. The total value of the money is $725.00. Find the number of each bill.
Please explain the steps in solving this problem.
Thank you

Answer by Alan3354(69443)

A bank teller has some five dollar bills and some twenty dollar bills. The teller has five more twenties than fives. The total value of the money is $725.00. Find the number of each bill.
--------------------
f = # of 5's
t = # of 20's
---------------
t = f + 5
5f + 20t = 725
Sub for t
5f + 20(f+5) = 725
5f + 20f+100 = 725
25f = 625
f = 25 $5 bills
t = 30 $20 bills